OpenStudy (anonymous):

HELP ASAP. MEDAL! solve the equation sqrty+8=11

geerky42 (geerky42):

Like this? \(\Large \sqrt{y+8}=11\)

OpenStudy (anonymous):

3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

no it would be 9

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

sorry i didnt see the sqroot

OpenStudy (anonymous):

the sqroot of y + 8 = 11 the sqroot of y is 3 cuz 3 + 8 is 11 to un sqroot something you multiply it by itself so it would b 9

geerky42 (geerky42):

@athenasannabeth It is correct if the problem is \(\sqrt y + 8 = 11\), but @baybay88 already confirmed it is \(\sqrt{y+8}=11\) In that case, you just square both sides to get \(y + 8 = 121 \rightarrow \boxed{y = 113}\)